OpenStudy (anonymous):

find k such that y=ksqrt(x) is tangent to y=x+6

OpenStudy (anonymous):

calculus is the best process to solve this question

OpenStudy (anonymous):

The procedure is very much like your last question.

OpenStudy (anonymous):

the derivative of the second equation is one. is the derivative of the first 1/2x?

OpenStudy (anonymous):

CALCULUS WAY\[m=\frac{ k }{ 2\sqrt{x} }\] \[\frac{ k }{ 2 \sqrt{x} }(1)=-1\] \[k=-2\sqrt{x}\]

OpenStudy (anonymous):

\[x+6=(-2\sqrt{x})\sqrt{x}\] \[x+6=-2x\] \[x=2,k \rightarrow -2 \sqrt{2}\]

OpenStudy (anonymous):

that's not the answer but thank u

OpenStudy (anonymous):

i think the question is supposed to be "find k such that y=x+6 is tangent to y=ksqrt(x) "